# Nonequilibrium Probes of Quantum Geometry in Gapless Systems

### v1

@@ 2.3 Quantum geometric tensor in CFT
${\cal G}_{m,n}$ $=\frac{1}{2}\langle h|\bigl\{L_{-m},L_{-n}\bigr\}|h\rangle-\langle h|L_{-m}|h\rangle\langle h|L_{-n}|h\rangle=|m|\Big(h+\frac{c}{24}(m^{2}-1)\Big)\delta_{m+n,0},$ (2.23a)
${\cal F}_{m,n}$ $\equiv-i\langle h|[L_{-m},L_{-n}]|h\rangle=2im\Big(h+\frac{c}{24}(m^{2}-1)\Big)\delta_{m+n,0}.$ (2.23b)
@@
That there are infinitely many such components reflects the fact that parameter space is infinite-dimensional. That all these components vanish when $m=0$ or $n=0$ reflects the fact that parameter space is a quotient (2.12). Both the metric and the curvature are nondegenerate on the manifold (2.12) if $h>0$; by contrast, if $h=0$, the directions $m=\pm 1$ become degenerate due to the vanishing factor $m^{2}-1$. This reflects the fact that the quantum geometry of coherent states (2.17) is that of a coadjoint orbit [59]—in the case at hand, an orbit of the Virasoro group. Finally, that the components (2.23) are identical up to an absolute value of $m\in\mathbb{Z}$ in the metric (and a trivial rescaling) reflects the fact that the quantum geometry on the manifold (2.12) is Kähler, as is indeed the case for Virasoro coadjoint orbits [57, 4]. This Kähler property is also true for Bloch coherent states and squeezed states, of which our current setup is an infinite-dimensional analogue.
@@
While it is convenient to write the metric and Berry curvature in terms of Fourier modes as in Eq. (2.22), it is also possible to express them as functionals of local vector fields. For the curvature, this can be done e.g. by inverse Fourier transforming Eq. (2.22b), which yields
@@
${\cal F}(\mathbf{X},\mathbf{Y})=-2\oint\frac{\textrm{d}\varphi}{2\pi}\Big[\Big(h-\frac{c}{24}\Big)XY^{\prime}-\frac{c}{24}XY^{\prime\prime\prime}\Big].$ (2.24)
@@
The analogous computation for the quantum metric is more involved, as the metric turns out to be a nonlocal functional of $\mathbf{X},\mathbf{Y}$. This is due to the absolute value of $m$ in Eq. (2.23a), which gives rise to the propagator $\sum_{m\in\mathbb{Z}}|m|e^{im(\theta-\varphi)}=-[2\sin^{2}([\theta-\varphi]/2)]^{-1}$. Thus, one eventually finds
${\cal G}(\mathbf{X},\mathbf{Y})=\oint\frac{\textrm{d}\varphi\,\textrm{d}\theta}{16\pi^{2}}\frac{(h{-}\frac{c}{24})[X(\varphi){-}X(\theta)][Y(\varphi){-}Y(\theta)]+\frac{c}{24}[X^{\prime}(\varphi){-}X^{\prime}(\theta)][Y^{\prime}(\varphi){-}Y^{\prime}(\theta)]}{\sin^{2}([\theta-\varphi]/2)}$ (2.25)
@@
for the Virasoro quantum metric at the identity. Note how different this is from the Berry curvature (2.24), all because of the seemingly innocuous distinction between (2.23a) and (2.23b). In this sense, the Kähler structure of the parameter space (2.12) is manifest in Fourier space, but not in position space.
3 Perturbative quantum geometry
@@
The simplest probes of quantum geometry for the coherent states (2.17) are provided by perturbative drives, i.e. deformed Hamiltonians (2.13) whose $f_{t}$ is infinitesimally close to the identity. Such setups cannot probe global properties (which involve finite $f_{t}$), but do suffice to determine the quantum geometric tensor at a single point in parameter space, namely that of the undeformed Hamiltonian (2.10). Then, the key simplification is that the Lie algebra (2.9) determines everything, without involving finite conformal transformations. Homogeneity of the parameter space (2.12) further means that quantum geometry at the identity determines quantum geometry at any other point, so there is no loss of generality in focusing on infinitesimal drives as far as local quantum geometry is concerned.
@@
In this section, we consider such perturbative drives to derive two universal relations: (i) the link between quantum metric and absorption rates, (ii) the link between Berry curvature and linear-response coefficients.
3.1 Virasoro quantum metric from absorption rates
Note from the quantum metric (2.21a) that the squared norm of a vector at the identity is
${\cal G}(\mathbf{X},\mathbf{X})=-\langle h|\mathfrak{u}(\mathbf{X})^{2}|h\rangle+\langle h|\mathfrak{u}(\mathbf{X})|h\rangle^{2}.$ (3.1)
@@ 3.1 Virasoro quantum metric from absorption rates
This is nothing but the variance of the Hermitian operator $i\mathfrak{u}(\mathbf{X})$. Keeping in mind the fluctuation-dissipation theorem, it suggests that the quantum metric can be probed with admittance experiments. Such a proposal was indeed put forward in [52], and we apply it here to CFT.
Transition rates from perturbative Virasoro drives.
@@
Consider a time-dependent Hamiltonian (2.13) whose deformation $f_{t}(\varphi)=\varphi+\epsilon X_{t}(\varphi)$ is close to the identity in the sense that $\epsilon$ is small, while $\mathbf{X}_{t}$ is some finite, time-dependent vector field that does not depend on $\epsilon$. Then, to first order in $\epsilon$, the perturbed Hamiltonian reads
$H(t)=H_{0}+\epsilon\bigl[\mathfrak{u}(\mathbf{X}_{t}),H_{0}\bigr]+O(\epsilon^{2}),$ (3.2)
@@
where $H_{0}$ is the homogeneous Hamiltonian (2.10) and $\mathfrak{u}(\cdot)$ is the algebra representation (2.6). We assume throughout that $X_{t}(\varphi)$ vanishes for $t<0$, i.e. the perturbation is ‘switched on’ at $t=0$. In the present case, we further assume that the perturbation for $t>0$ is harmonic with frequency $\omega>0$ so that
@@
$X_{t}(\varphi)=\begin{cases}0&\text{for}\;\,t<0,\vskip 4.2679pt\\ \displaystyle\frac{1}{\omega L}\cos(\omega t)X(\varphi)&\text{for}\;\,t>0,\end{cases}$ (3.3)
@@
The Hamiltonian (3.2) with $\mathbf{X}_{t}$ given by Eq. (3.3) falls into the usual regime of time-dependent perturbation theory. In particular, one can perturbatively compute late-time transition rates between eigenstates of $H_{0}$. We focus here on transitions from the highest-weight state $|h\rangle$ of Eq. (2.16) to any other eigenstate, say $|\phi\rangle$, with an energy difference $E_{\phi}>0$ above that of $|h\rangle$. To leading order in $\epsilon$, the resulting transition rate is
$\Gamma_{\phi}(\omega)\sim\frac{\epsilon^{2}}{2\omega^{2}L^{2}}\Bigl|\big\langle\phi\big|\bigl[\mathfrak{u}(\mathbf{X}),H_{0}\bigr]\big|h\big\rangle\Bigr|^{2}\delta\big(\omega-E_{\phi}\big)=\frac{\epsilon^{2}}{2L^{2}}\bigl|\langle\phi|\mathfrak{u}(\mathbf{X})|h\rangle\bigr|^{2}\delta\big(\omega-E_{\phi}\big).$ (3.4)
@@
Integrated absorption rate and quantum metric.
Eq. (3.4) can be used to obtain the absorption rate induced by the perturbation (3.2) by summing over all possible final states of the transition. Since the deformed Hamiltonian (3.2) is an infinitesimal form of unitary conjugation (2.11), the only nonzero transition amplitudes are those whose final state $|\phi\rangle$ is a descendant state of $|h\rangle$, i.e. one that belongs to the space of the representation based on $|h\rangle$. The absorption rate thus reads
$\Gamma(\omega)\sim\frac{\epsilon^{2}}{2L^{2}}\sum_{k=1}^{\infty}\bigl|\langle\phi_{k}|\mathfrak{u}(\mathbf{X})|h\rangle\bigr|^{2}\delta(\omega-E_{\phi_{k}}),$ (3.5)
where we label each descendant of $|h\rangle$ by an integer $k=1,2,\ldots$. How exactly one chooses the labeling has no importance: the only key points are (i) that the spectrum of $H_{0}$ is discrete thanks to the finite size $L$, and (ii) that $|h\rangle$ is a nondegenerate eigenstate of $H_{0}$ within its irreducible representation.
Integrating the rate (3.5) over all frequencies yields
$\int_{0}^{\infty}\textrm{d}\omega\,\Gamma(\omega)\sim\frac{\epsilon^{2}}{2L^{2}}\sum_{k=1}^{\infty}\bigl|\langle\phi_{k}|\mathfrak{u}(\mathbf{X})|h\rangle\bigr|^{2}=\frac{\epsilon^{2}}{2L^{2}}\Bigl(\langle h|\mathfrak{u}(\mathbf{X})|h\rangle^{2}-\langle h|\mathfrak{u}(\mathbf{X})^{2}|h\rangle\Bigr)$ (3.6)
upon using that $\mathfrak{u}(\cdot)$ is anti-Hermitian and writing the identity operator as $\sum_{k=1}^{\infty}|\phi_{k}\rangle\langle\phi_{k}|+|h\rangle\langle h|$ within the irreducible representation of $|h\rangle$. The result (3.6) manifestly involves the variance of $i\mathfrak{u}(\mathbf{X})$, which we found in Eq. (3.1) to be the squared quantum norm of $\mathbf{X}$. In other words,
$\int_{0}^{\infty}\textrm{d}\omega\,\Gamma(\omega)\sim\frac{\epsilon^{2}}{2L^{2}}\,{\cal G}(\mathbf{X},\mathbf{X})$ (3.7)
where $\cal G$} is the quantum metric at the identity coset in the infinite-dimensional parameter space (2.12). This is the result announced in Eq. (1.4). It is consistent with the protocol proposed earlier in [52], applied there to two-level systems and Bloch bands. As in [52], the same protocol can be used to measure any value ${\cal G}(\mathbf{X},\mathbf{Y})$, since ${\cal G}(\mathbf{X},\mathbf{Y})=[{\cal G}(\mathbf{X}+\mathbf{Y},\mathbf{X}+\mathbf{Y})-{\cal G}(\mathbf{X},\mathbf{X})-{\cal G}(\mathbf{Y},\mathbf{Y})]/2$.
3.2 Virasoro Berry curvature from linear response
Let us now link the Berry curvature (2.21b) to the linear response of the stress-energy tensor to perturbations of the form (3.2). This serves several purposes: it highlights the Kähler structure of the parameter space (2.12), it reaffirms the ballistic transport properties of CFT [24] in the language of viscosity, and it lays the foundations for the adiabatic considerations of Sec. 4. We will also briefly return to linear and nonlinear responses in Sec. 5.
@@ 3.2 Virasoro Berry curvature from linear response
Consider again a time-dependent Hamiltonian $H(t)$ of the form (3.2), for a system prepared in the eigenstate $|h\rangle$ of $H_{0}$ at $t=0$. In contrast to Sec. 3.1, let the vector field $\mathbf{X}_{t}=X_{t}(\varphi)\partial_{\varphi}$ have an arbitrary time-dependence, with modes $\hat{X}_{m}(t)$ in the Fourier decomposition (2.5). We are interested in the ensuing response of expectation values of Hermitian operators $i\mathfrak{u}(\mathbf{Y})$ generated by the stress-energy tensor (2.6) for any vector field $\mathbf{Y}=Y(\varphi)\partial_{\varphi}$. Denote by $|\phi(t)\rangle$ the state at time $t$ governed by the full time-dependent Hamiltonian. Then, standard linear-response arguments in the spirit of [60] yield (see Appendix C)
$\langle\phi(t)|i\mathfrak{u}(\mathbf{Y})|\phi(t)\rangle=\hat{Y}_{0}\Bigl(h-\frac{c}{24}\Bigr)+\epsilon\sum_{m,n\neq 0}\hat{Y}_{m}\int_{0}^{t}\textrm{d}s\,\chi_{m,n}(t,s)\hat{X}_{n}(s)+O(\epsilon^{2})$ (3.8)
@@
$\chi_{m,n}(t,s)\equiv\partial_{s}{\cal F}_{m,n}e^{i2\pi(mt+ns)/L}=-i\frac{2\pi}{L}m{\cal F}_{m,-m}e^{i2\pi m(t-s)/L}\delta_{m+n,0}$ (3.9)
in terms of the Berry curvature (2.23b) in Fourier space. Formally picking $\mathbf{Y}=\delta(\varphi-\theta)\partial_{\varphi}$, i.e. the modes $\hat{Y}_{m}=e^{-im\theta}/{2\pi}$, one immediately finds
$\langle\phi(t)|{\mathcal{T}}(\theta)|\phi(t)\rangle=\frac{1}{2\pi}\Bigl(h-\frac{c}{24}\Bigr)-\epsilon\int_{0}^{t}\frac{\textrm{d}s}{L}\Bigl[\Bigl(h-\frac{c}{24}\Bigr)\partial_{\theta}^{2}-\frac{c}{24}\partial_{\theta}^{4}\Bigr]X_{s}(\theta-t+s)+O(\epsilon^{2})$ (3.10)
for the stress-energy tensor (2.7). This amounts to rewriting Eqs. (3.8)–(3.9) in terms of the real-space, functional Berry curvature (2.24) rather than in Fourier space. Note that the zero mode of $X_{t}(\varphi)$ does not contribute, as was to be expected on the parameter space (2.12).
@@
The perturbation $X_{t}(\varphi)$ is ‘adiabatic’ when each Fourier mode $\hat{X}_{m}(t)$ (with $m\neq 0$ without loss of generality) satisfies $L|\partial_{t}^{k+1}\hat{X}_{m}(t)|\ll|\partial_{t}^{k}\hat{X}_{m}(t)|$ for any $k\geq 1$. In this regime, it is meaningful to expand expectation values in time derivatives of the perturbation. Eq. (3.8) then becomes (again, see Appendix C)
@@
$\langle\phi(t)|i\mathfrak{u}(\mathbf{Y})|\phi(t)\rangle\sim\hat{Y}_{0}\Bigl(h-\frac{c}{24}\Bigr)+\epsilon\sum_{m\neq 0}{\cal F}_{m,-m}\hat{Y}_{m}\biggl(1+\frac{L\partial_{t}}{2\pi im}+O\bigl(L^{2}\partial_{t}^{2}/m^{2}\bigr)\biggr)\hat{X}_{-m}(t)$ (3.11)
to linear order in $\epsilon$, again involving ${\cal F}_{m,n}$ in Eq. (2.23b). The adiabatic version of Eq. (3.10) is
@@
$\langle\phi(t)|{\mathcal{T}}(\varphi)|\phi(t)\rangle\sim\frac{1}{2\pi}\Bigl(h-\frac{c}{24}\Bigr)-\frac{\epsilon}{\pi}\Bigl(h-\frac{c}{24}-\frac{c}{24}\partial_{-}^{2}\Bigr)\partial_{-}\bigl[X_{t}(\varphi)-\hat{X}_{0}(t)\bigr]$ (3.12)
@@
in terms of the partial derivative $\partial_{-}\equiv\partial_{\varphi}-\tfrac{L}{2\pi}\partial_{t}$ along the light-cone coordinate $\varphi^{-}$ defined above Eq. (2.1).
@@
The links to quantum geometry are manifest, as all formulas since Eq. (3.8) involve the Berry curvature (2.23b) or its functional analogue (2.24). In the case at hand, Berry curvature and Virasoro algebra are so closely related that one can nearly read off the Virasoro commutation relations from the linear response, e.g. through Eq. (3.12) with the ubiquitous higher-derivative contribution of the central charge $c$. (Similar formulas were previously obtained in [26] for inhomogeneous CFT out of equilibrium.) Note also the similarity between Eqs. (3.11)–(3.12) and the standard definition of viscosity, where the stress-energy tensor depends on the rate of strain. Indeed, the odd viscosity of quantum Hall systems is precisely their linear response to adiabatic deformations of the metric [47]. Such deformations are implemented by unitary operators [48, 51], similar to what we did in Eq. (2.11) for CFTs. We stress, however, that the viscosity tensor vanishes in standard 1+1D CFT: its odd part is trivially zero for dimensional reasons, and its dissipative part vanishes since all perturbations propagate ballistically. Instead, our result (3.12) should be seen as a response of the thermal current to perturbations of the metric, i.e. a ballistic contribution to the thermal conductance of a 1+1D CFT [26, 24].
4 Quantum geometry from adiabatic drives

### v2

@@ II.3 Quantum geometric tensor
${\cal G}_{m,n}$ $\equiv\frac{1}{2}\langle h|\bigl\{L_{-m},L_{-n}\bigr\}|h\rangle-\langle h|L_{-m}|h\rangle\langle h|L_{-n}|h\rangle$
$=|m|\Big(h+\frac{c}{24}(m^{2}-1)\Big)\delta_{m+n,0},$ (30a)
${\cal F}_{m,n}$ $\equiv-i\langle h|[L_{-m},L_{-n}]|h\rangle$
$=2im\Big(h+\frac{c}{24}(m^{2}-1)\Big)\delta_{m+n,0}.$ (30b)
@@
That there are infinitely many components reflects the fact that parameter space is infinite-dimensional. The components with $m=0$ or $n=0$ vanish, because parameter space is a quotient (18). Both the metric and the curvature are nondegenerate on the manifold (18) if $h>0$; by contrast, if $h=0$, the directions $m=\pm 1$ become degenerate due to the vanishing factor $m^{2}-1$ in Eq. (30). [^11] The cancellation stems from the fact that the CFT vacuum $|0\rangle$ defined below Eq. (22) is invariant under the $\mathrm{SL}(2,\mathbb{R})$ group generated by $L_{-1}$, $L_{0}$, $L_{1}$, which is not true for a generic highest-weight state $|h\rangle$. We return to this in Sec. V. This is ultimately because the quantum geometry of coherent states (23) is that of a coadjoint orbit [86]—in the case at hand, an orbit of the Virasoro group [87, 88, 57, 89]. Finally, that the components (30) are identical up to an absolute value of $m\in\mathbb{Z}$ in the metric (and a trivial rescaling) reflects the fact that the quantum geometry on the manifold (18) is Kähler, as is indeed the case for Virasoro coadjoint orbits [56, 57, 3]. One specifically has ${\cal G}(\mathbf{X},\mathbf{Y})=\tfrac{1}{2}{\cal F}(\mathbf{X},J\mathbf{Y})$, where the complex structure $J$ maps each Fourier mode $\hat{Y}_{n}$ on $i\operatorname{sgn}(n)\hat{Y}_{n}$. This Kähler property is also true for Bloch coherent states and squeezed states, of which our current setup is an infinite-dimensional analogue.
@@
While it is convenient to write the metric and Berry curvature in terms of Fourier modes as in Eq. (29), it is also possible to express them as functionals of local vector fields. For the curvature, this can be done e.g. by inverse Fourier transforming Eq. (29b), which yields
@@
${\cal F}(\mathbf{X},\mathbf{Y})=-2\oint\frac{\textrm{d}\varphi}{2\pi}\Bigl[\Bigl(h-\frac{c}{24}\Bigr)XY^{\prime}-\frac{c}{24}XY^{\prime\prime\prime}\Bigr].$ (31)
@@
The analogous computation for the quantum metric is more involved, as the metric turns out to be a nonlocal functional of $\mathbf{X},\mathbf{Y}$. This is due to the absolute value of $m$ in Eq. (30a), which gives rise to the propagator $\sum_{m\in\mathbb{Z}}|m|e^{im(\theta-\varphi)}=-[2\sin^{2}([\theta-\varphi]/2)]^{-1}$. Thus, one eventually finds
${\cal G}(\mathbf{X},\mathbf{Y})$ $=\oint\frac{\textrm{d}\varphi\,\textrm{d}\theta}{16\pi^{2}\sin^{2}([\varphi-\theta]/2)}$ (32)
$\quad\times\biggl\{\Bigl(h-\frac{c}{24}\Bigr)[X(\varphi){-}X(\theta)][Y(\varphi){-}Y(\theta)]$
$\quad\qquad+\frac{c}{24}[X^{\prime}(\varphi){-}X^{\prime}(\theta)][Y^{\prime}(\varphi){-}Y^{\prime}(\theta)]\biggr\}$
@@
for the Virasoro quantum metric at the identity. Note how different this is from the Berry curvature (31), all because of the seemingly innocuous distinction between (30a) and (30b). In this sense, the Kähler structure of the parameter space (18) is manifest in Fourier space, but not in position space.
III Perturbative quantum geometry
@@
The simplest probes of quantum geometry for the coherent states (23) are provided by perturbative drives, i.e. deformed Hamiltonians (19) whose $f_{t}$ is infinitesimally close to the identity. Such setups cannot probe global properties (which involve finite $f_{t}$), but they suffice to determine the quantum geometric tensor at a single point in parameter space, namely that of the undeformed Hamiltonian (16). Then, the key simplification is that the Lie algebra (15) determines everything, without involving finite conformal transformations. Homogeneity of the parameter space (18) further means that quantum geometry at the identity determines quantum geometry at any other point, so there is no loss of generality to focus on infinitesimal drives as far as local quantum geometry is concerned.
@@
In this section, we consider such perturbative drives to derive two universal relations between (i) the quantum metric and absorption rates, and (ii) the Berry curvature and linear responses.
III.1 Quantum metric from absorption rates
@@
Keeping in mind the fluctuation-dissipation theorem, the expression (28) for the squared quantum norm as a variance suggests that the quantum metric can be probed in admittance experiments [74, 75, 76, 77]. Such a proposal was indeed put forward in [73]. Here, we develop these ideas for the infinite-dimensional quantum geometry of 1+1D CFT.
Transition rates and perturbative drives.
@@
Consider a time-dependent Hamiltonian (19) whose deformation $f_{t}(\varphi)=\varphi+\epsilon X_{t}(\varphi)$ is close to the identity in the sense that $\epsilon$ is small, while $\mathbf{X}_{t}$ is some unspecified time-dependent vector field. Then, to first order in $\epsilon$, the perturbed Hamiltonian reads
$H(t)=H_{0}+\epsilon\bigl[\mathfrak{u}(\mathbf{X}_{t}),H_{0}\bigr]+O(\epsilon^{2}),$ (33)
@@
where $H_{0}$ is the homogeneous Hamiltonian (16) and $\mathfrak{u}(\cdot)$ is the algebra representation (12). We assume throughout that $X_{t}(\varphi)$ vanishes for $t<0$, i.e. the perturbation is ‘switched on’ at $t=0$. In the present case, we further assume that the perturbation for $t>0$ is harmonic with frequency $\omega>0$ so that
@@
$X_{t}(\varphi)=\left\{\begin{aligned} &0&&\text{for}\;\,t<0,\vskip 4.2679pt\\ &\displaystyle\frac{1}{\omega L}\cos(\omega t)X(\varphi)&&\text{for}\;\,t>0,\end{aligned}\right$. (34)
@@
The Hamiltonian (33) for $\mathbf{X}_{t}$ given by Eq. (34) falls into the usual regime of time-dependent perturbation theory. In particular, one can perturbatively compute late-time transition rates between eigenstates of $H_{0}$. We focus here on transitions from the highest-weight state $|h\rangle$ of Eq. (22) to any other eigenstate, say $|\phi\rangle$, with an energy difference $E_{\phi}>0$ above that of $|h\rangle$. To leading order in $\epsilon$, the resulting transition rate is
$\Gamma_{\phi}(\omega)$ $\sim\frac{\epsilon^{2}}{2\omega^{2}L^{2}}\Bigl|\big\langle\phi\big|\bigl[\mathfrak{u}(\mathbf{X}),H_{0}\bigr]\big|h\big\rangle\Bigr|^{2}\delta\big(\omega-E_{\phi}\big)$ (35)
$=\frac{\epsilon^{2}}{2L^{2}}\bigl|\langle\phi|\mathfrak{u}(\mathbf{X})|h\rangle\bigr|^{2}\delta\big(\omega-E_{\phi}\big).$
@@
Integrated absorption rates.
Eq. (35) can be used to obtain the absorption rate induced by the perturbation (33) by summing over all possible final states of the transition. Since the deformed Hamiltonian (33) is an infinitesimal form of unitary conjugation (17), the only nonzero transition amplitudes are those whose final state $|\phi\rangle$ is a descendant state of $|h\rangle$, i.e. one that belongs to the space of the representation based on $|h\rangle$. The absorption rate thus reads
$\Gamma(\omega)\sim\frac{\epsilon^{2}}{2L^{2}}\sum_{k=1}^{\infty}\bigl|\langle\phi_{k}|\mathfrak{u}(\mathbf{X})|h\rangle\bigr|^{2}\delta(\omega-E_{\phi_{k}}),$ (36)
where we label each descendant $|\phi_{k}\rangle$ of $|h\rangle$ by an integer $k=1,2,\ldots$. How exactly one chooses the labeling has no importance: the only key points are (i) that the spectrum of $H_{0}$ is discrete thanks to the finite size $L$, and (ii) that $|h\rangle$ is a nondegenerate eigenstate of $H_{0}$ within its irreducible representation.
Integrating the rate (36) over all frequencies yields
$\int_{0}^{\infty}\textrm{d}\omega\,\Gamma(\omega)$ $\sim\frac{\epsilon^{2}}{2L^{2}}\sum_{k=1}^{\infty}\bigl|\langle\phi_{k}|\mathfrak{u}(\mathbf{X})|h\rangle\bigr|^{2}$ (37)
$=\frac{\epsilon^{2}}{2L^{2}}\Bigl(\langle h|\mathfrak{u}(\mathbf{X})|h\rangle^{2}-\langle h|\mathfrak{u}(\mathbf{X})^{2}|h\rangle\Bigr),$
upon using that $\mathfrak{u}(\cdot)$ is anti-Hermitian and writing the identity operator as $\sum_{k=1}^{\infty}|\phi_{k}\rangle\langle\phi_{k}|+|h\rangle\langle h|$ within the irreducible representation of $|h\rangle$. The result (37) manifestly involves the variance of $i\mathfrak{u}(\mathbf{X})$, which we found in Eq. (28) to be the squared quantum norm of $\mathbf{X}$. In other words,
$\int_{0}^{\infty}\textrm{d}\omega\,\Gamma(\omega)\sim\frac{\epsilon^{2}}{2L^{2}}\,{\cal G}(\mathbf{X},\mathbf{X}),$ (38)
where $\cal G$} is the quantum metric at the identity coset in the infinite-dimensional parameter space (18). This is the result announced in Eq. (4). It is consistent with the protocol proposed earlier in [73], applied there to two-level systems and Bloch bands. As in [73], the same protocol can be used to obtain any value ${\cal G}(\mathbf{X},\mathbf{Y})$, since ${\cal G}(\mathbf{X},\mathbf{Y})=[{\cal G}(\mathbf{X}+\mathbf{Y},\mathbf{X}+\mathbf{Y})-{\cal G}(\mathbf{X},\mathbf{X})-{\cal G}(\mathbf{Y},\mathbf{Y})]/2$.
III.2 Berry curvature from linear response
Let us now link the Berry curvature (27b) to the linear response of the stress-energy tensor to perturbations of the form (33). This serves several purposes: it highlights the Kähler structure of the parameter space (18), it reaffirms the ballistic transport properties of CFT [33] in the language of viscosity, and it lays the foundations for the adiabatic considerations of Sec. IV, where Virasoro Berry phases [46] will make an appearance. We will also briefly return to linear and nonlinear responses in Sec. V.
@@ III.2 Berry curvature from linear response
Consider again a time-dependent Hamiltonian $H(t)$ of the form (33), for a system prepared in the eigenstate $|h\rangle$ of $H_{0}$ at $t=0$. In contrast to Sec. III.1, let the vector field $\mathbf{X}_{t}=X_{t}(\varphi)\partial_{\varphi}$ have an arbitrary time-dependence, with modes $\hat{X}_{m}(t)$ in the Fourier decomposition (11). We are interested in the ensuing response of expectation values of Hermitian operators $i\mathfrak{u}(\mathbf{W})$ generated by the stress-energy tensor (12) for any vector field $\mathbf{W}=W(\varphi)\partial_{\varphi}$. Denote by $|\phi(t)\rangle$ the state at time $t$ governed by the full time-dependent Hamiltonian. Then, standard linear-response arguments in the spirit of [90] yield (see Appendix C)
$\langle\phi(t)|i\mathfrak{u}(\mathbf{W})|\phi(t)\rangle=\hat{W}_{0}\Bigl(h-\frac{c}{24}\Bigr)$ (39)
$\qquad+\epsilon\sum_{m,n\neq 0}\hat{W}_{m}\int_{0}^{t}\textrm{d}s\,\chi_{m,n}(t,s)\,\hat{X}_{n}(s)+O(\epsilon^{2})$
@@
$\chi_{m,n}(t,s)$ $\equiv\partial_{s}{\cal F}_{m,n}e^{2\pi i(mt+ns)/L}$ (40)
$=-i\frac{2\pi}{L}m{\cal F}_{m,-m}e^{2\pi im(t-s)/L}\delta_{m+n,0}$
in terms of the Berry curvature (29b)–(30b) in Fourier space. This is the result announced in Eq. (5), written here in Fourier modes, but expressible in real space using Eq. (31). Formally picking $\mathbf{W}=\delta(\varphi-\theta)\partial_{\varphi}$, i.e. the modes $\hat{W}_{m}=e^{-im\theta}/{2\pi}$, one immediately finds
$\langle\phi(t)|{\mathcal{T}}(\theta)|\phi(t)\rangle$ (41)
$=\frac{1}{2\pi}\Bigl(h-\frac{c}{24}\Bigr)-2\epsilon\int_{0}^{t}\frac{\textrm{d}s}{L}\Bigl[\Bigl(h-\frac{c}{24}\Bigr)\partial_{\theta}^{2}-\frac{c}{24}\partial_{\theta}^{4}\Bigr]$
$\qquad\qquad\qquad\qquad\quad\times X_{s}\Bigl(\theta-2\pi[t-s]/L\Bigr)+O(\epsilon^{2})$
for the stress-energy tensor. This amounts to rewriting Eqs. (39)–(40) in terms of the real-space, functional Berry curvature (31) rather than in Fourier space. Note that the zero mode of $X_{t}(\varphi)$ does not contribute, as was to be expected on the parameter space (18). Lastly, while we restricted ourselves here to linear response, one may similarly compute nonlinear-response terms, which probe higher-order quantum-geometric quantities [91].
@@
The perturbation $X_{t}(\varphi)$ is ‘adiabatic’ when each Fourier mode $\hat{X}_{m}(t)$ (with $m\neq 0$ without loss of generality) satisfies $L|\partial_{t}^{k+1}\hat{X}_{m}(t)|\ll|\partial_{t}^{k}\hat{X}_{m}(t)|$ for any $k\geq 1$. In this regime, it is meaningful to expand expectation values in time derivatives of the perturbation. Eq. (39) then becomes (again, see Appendix C)
@@
$\langle\phi(t)|i\mathfrak{u}(\mathbf{W})|\phi(t)\rangle\sim\hat{W}_{0}\Bigl(h-\frac{c}{24}\Bigr)+\epsilon\sum_{m\neq 0}{\cal F}_{m,-m}$
$\times\hat{W}_{m}\biggl(1+\frac{L\partial_{t}}{2\pi im}+O\bigl(L^{2}\partial_{t}^{2}/m^{2}\bigr)\biggr)\hat{X}_{-m}(t)$ (42)
to first order in $\epsilon$, again involving ${\cal F}_{m,n}$ in Eq. (30b). The adiabatic version of Eq. (41) is
$\langle\phi(t)|{\mathcal{T}}(\varphi)|\phi(t)\rangle\sim\frac{1}{2\pi}\Bigl(h-\frac{c}{24}\Bigr)$ (43)
@@
$\qquad\quad-\frac{\epsilon}{\pi}\Bigl(h-\frac{c}{24}-\frac{c}{24}\partial_{-}^{2}\Bigr)\partial_{-}\bigl[X_{t}(\varphi)-\hat{X}_{0}(t)\bigr]$
@@
in terms of the partial derivative $\partial_{-}\equiv\partial_{\varphi}-\tfrac{L}{2\pi}\partial_{t}$ along the light-cone coordinate $\varphi^{-}$ defined above Eq. (7).
@@
The links to quantum geometry are manifest, as all formulas since Eq. (39) involve the Berry curvature (30b) or its functional analogue (31). In the case at hand, Berry curvature and Virasoro algebra are so closely related that one can nearly read off the Virasoro commutation relations from the linear response, e.g. through Eq. (43) with the ubiquitous higher-derivative contribution proportional to the central charge $c$. (Similar formulas were previously obtained in [35] for inhomogeneous CFT out of equilibrium.) Note also the similarity between Eqs. (III.2)–(43) and the standard definition of viscosity, where the stress-energy expectation value depends on the rate of strain. Indeed, the odd viscosity of quantum Hall systems is precisely their linear response to adiabatic deformations of the metric [68]. Such deformations are implemented by unitary operators [69, 72], similar to our Eq. (17) for CFTs. We stress, however, that the viscosity tensor vanishes in standard 1+1D CFT: its odd part is trivially zero for dimensional reasons, and its dissipative part vanishes since all perturbations propagate ballistically. Instead, our result (43) should be seen as a response of the thermal current to perturbations of the metric, i.e. a ballistic contribution to the thermal conductance of a 1+1D CFT [35, 33].
IV Adiabatic quantum geometry